### v2

# Visual and auditory insights into multifrequency variability

Abstract
Using open astronomical multifrequency databases, we constructed light curves and developed a comprehensive visualisation and sonification analysis for the blazars Mrk 501, Mrk 1501, Mrk 421, BL Lacerta, AO 0235+164, 3C 66A, OJ 049, OJ 287, and PKS J2134-0153. This study employed Musical Instrument Digital Interface (MIDI) and Parameter Mapping Sonification (PMSon) techniques to generate waveforms, spectrograms, and sonifications. These representations demonstrate that data visualisation and sonification are powerful tools for analysing astronomical objects like blazars, providing insights into their multifrequency variability. This work highlights how sonification and visualisation can aid in identifying potential patterns, power variations, regularities, and gaps in the data. This multimodal approach also underscores the importance of inclusivity in scientific communication, offering accessible methods for exploring the complex behaviour of blazars.
keywords:
astronomical data bases: miscellaneous, virtual observatory tools, software: data analysis
pubyear: 2025pagerange: Exploring blazars through sonification. Visual and auditory insights into multifrequency variabilityExploring blazars through sonification. Visual and auditory insights into multifrequency variability
1 Introduction
Data visualisation, often referred to as DataViz, serves as a valuable approach for identifying and discerning patterns or recurring trends within the behaviours of social or natural phenomena. It leverages its cognitive and explanatory capabilities through techniques such as time series, histograms, infographics, and Venn-Euler diagrams (Friendly, 2008). Consequently, graphical representation stands as a fundamental method for conducting research in the realms of science and study.
DataViz stands as a important methodology for both scientific communication and research. Nonetheless, it is essential to consider a couple of factors. Firstly, it is important to recognise that visual communication is not accessible to individuals who are blind or have visual impairments (BVI) (Pérez-Montero, 2019). Secondly, relying solely on graphical representations might prove insufficient when dealing with comprehensive data analysis. Therefore, incorporating supplementary review methods becomes indispensable to enhance the research process.
Taking this into account, data sonification can serve not only as a means of inclusivity in scientific communication but also as an additional tool for research (Sawe et al., 2020). By transforming data into sound, it offers an alternative mode of perception that complements visual representations, allowing for a deeper understanding of complex data.
As Kramer et al. (1999) defines it, data sonification as “the transformation of data relations into perceived relations in an acoustic signal for the purposes of facilitating communication or interpretation” highlights sonification as a valuable supplementary instrument for comprehending information. Unlike the sense of vision, hearing excels in perceiving information over time, as well as recognising patterns and momentary fluctuations. Therefore, sonification proves to be an excellent aid in enhancing knowledge understanding (Guttman et al., 2005).
Another definition of data sonification, provided by Hermann (2008), characterises it as a technique that takes data as its input and produces sound signals, possibly in response to optional additional stimulation or triggering. This broad definition encompasses any process that uses data to create auditory outputs and can be referred to as sonification, if and only if the following conditions are satisfied:
  • The sound reflects objective properties or relations of the input data.
  • The transformation is systematic. This means that there is a precise definition provided of how the data (and optional interactions) cause the sound to change.
  • The sonification is reproducible: given the same data and identical interactions (or triggers), the resulting sound has to be structurally identical.
  • The system can intentionally be used with different data and also be used in repetition with the same data.
Following Hermann (2008) and Vogt (2008), the “sonification is the data-dependent generation of sound, if the transformation is systematic, objective and reproducible, so that it can be used as scientific method.”
Numerous instances demonstrate the effectiveness of data sonification as a valuable science communication tool. Notable examples include its application in depicting salmon migration patterns (Hegg et al., 2018) and visualising fluctuations in brainwave activity (Parvizi et al., 2018). The Geiger counter represents yet another compelling case in point.
In the realm of astronomy, early instances of data sonification trace to Jansky (1933), when detected radio waves emanating from the centre of the Milky Way were analysed by telephone communication noise. Three decades later, Penzias & Wilson (1965) identified the interference in their observations as originating from cosmic microwave background radiation, marking a significant discovery in the field (Kellermann et al., 2020).
Another example is that of Morgan et al. (1997) who took the data of the X-ray emission of the black hole GRS 1915+105 and translated it into audio signals, allowing us to “hear” the accretion disk from it Masetti (2013). Also Chandra (2003) has sonified data and released “Hears” of a black hole for the first time.
Abbott et al. (2016) documented the simultaneous detection of a brief signal by two gravitational wave detectors, which resulted from the fusion of two black holes. The researchers converted the data from the instant of the collision, which they dubbed a “chirp.” This term was chosen because certain events producing gravitational waves exhibit auditory similarities to the chirping sounds of birds (LIGO Collaboration, 2016, see https://youtu.be/QyDcTbR-kEA).
Furthermore, Berti (2016) published an article entitled The First Sounds of Merging Black Holes, shedding light on the initial auditory cues of black hole mergers. In recent times, the National Aeronautics and Space Administration (NASA) has unveiled sonification imagery derived from optical data captured by the Hubble Space Telescope (NASA, 2022) and X-ray data observed by the Chandra X-ray Observatory (CXO) (Chandra, 2020).
In the context of communicating science to the BVI community, especially in the realm of astrophysics and black holes (BHs) this work sonifies data from nine blazars: Mrk 501, Mrk 1501, Mrk 421, BL Lacerta, AO 0235+164, 3C 66A (PKS 0219+428), OJ 049 (PKS 0829+046), OJ 287, and PKS J2134-0153.
[…]
With this background, in Section 2 we show the way to obtain the multifrequency data (radio, optical, X-ray, and $\gamma$-ray). Section 3 shows the sonification procedure: pre-production, production, and post-production. The discussion is shown in Section 4 with which we conclude our study.
2 Obtaining the data
2.1 Radio
The radio datasets for this work were obtained from two data bases: the University of Michigan Radio Astronomy Observatory (UMRAO) [^1] https://dept.astro.lsa.umich.edu/datasets/umrao.php, and the Astrogeo Very-long-baseline interferometry (Astrogeo-VLBI).
The 26-meter paraboloid UMRAO monitoring in the bands: 4.8, 8, and 14.5 GHz (Aller et al., 1985). It is focused on the study of total flux density and linear polarisation of active extragalactic objects [^2] https://lsa.umich.edu/astro/facilities/data-sets.html.
Radio Fundamental Catalog (RFC), of the Astro-geo Center, maintained by L. Petrov333 http://astrogeo.org/rfc provides precise positions with milli-arcseconds accuracies, maps, and estimates of correlated flux densities at baselines 1000–8000 km of thousands of compact radio sources.
2.2 Optical
For optical, the databases used were: American Association of Variable Star Observers (AAVSO) [^4] https://www.aavso.org/ and Zwicky Transient Facility (ZTF) [^5] https://www.ztf.caltech.edu/.
The optical AAVSO is a public database founded in 1910 and offers observations of variable stars collected and archived for world-wide access in collaboration with amateur and professional astronomers. It is an international organisation of long-term variable star observers who participate in scientific discovery through variable star astronomy.
[…]
2.3 X-rays
For the X-rays light curve, we use the Swift database with energies in the range $0.3-10\,\text{keV}$. It contains the X-Ray Telescope (XRT) with two important characteristics that makes it important for observations: low background and a constant point spread function across the field of view (Moretti et al., 2005).
2.4 Gamma-rays
The public database of $\gamma$-rays used was of The Fermi Large Area Telescope (LAT). This instrument observes the sky for $\sim$ 3 hours and covers $\gtrsim$ 20% of it for each observation. It has a crystal calorimeter for energy measurement, an anti-coincidence detector to distinguish the background of charged particles and high-resolution converter tracker for direction measurement of the incident gamma rays (Atwood et al., 2009). The $\gamma$-rays fluxes collected are in the range $100\,\text{MeV}-300\text{GeV}$.
3 Sonification
Sonifying the data for each of the blazar frequencies was carried out in three stages: pre-production, production and post-production.
3.1 Pre-production
Sonification pre-production basically focuses on data pre-processing. In the case of the radio and optical bands, the data was processed in such a way that any anomaly, disturbance or irrelevance of the data could be detected, eliminated, corrected or transformed. For this task the AWK pattern scanning and processing language was used.
In the case of X-rays, it was necessary to reduce the data and for this, the HEASoft [^6] https://heasarc.gsfc.nasa.gov/docs/software/lheasoft/ software was used. The data processing corresponding to the photons, the position and the orientation of the satellite is available on Flexible Image Transport System format files (FITS).
For the pre-production of sonification in the $\gamma$-rays, the data reduction was possible with the Fermitools [^7] https://fermi.gsfc.nasa.gov/ssc/data/analysis/software/ software. For this, it was necessary to know the main relevant parameters of the blazars such as: R.A., Dec., Region of Interest (ROI) and start and end time of the data processing (see e.g Magallanes-Guijón, 2020). The process of this reduction followed the procedure detailed by Cabrera et al. (2013).
[…]
3.1.1 Light curves
Each blazar selected had the following specifications:
(1) Mrk 501, has a redshift $z$ = 0.034 that corresponds to a distance $\sim$ 140 Mpc and R. A. = 16h 53m 52.2s, Dec. = +39° 45’ 37”. It has been monitored in radio (Richards et al., 2011), optical (Smith et al., 2009), X-rays (Abdo et al., 2011), and $\gamma$-rays (Dorner et al., 2017). This object has several periodicities reported (Bhatta, 2019). Recently Magallanes-Guijón & Mendoza (2024) have reported a periodicity $\sim$ 224 days.
The light curve in optical was built using the database from 1998 June 24 to 2021 September 12. The Swift database is from 2008 October 2 to 2021 October 8. And the $\gamma$-rays, covers a time interval from 2008 August 04 up to 2020 March 13. Table 1 shows specific information about the records and the light curve is presented in Figure 1.
Band Records Interim Total
  [y, m, d] [d]
Optical $11,849$ $23,2,17$ $8441$
X-rays $28,000$ $12,7,9$ $4607$
$\gamma$-rays $4,199$ $11,7,9$ $4239$
Table 1: The table shows each electromagnetic band studied for Mrk 501: the number of records, the interim in years, months, and days [y, m, d], and the total days for which at least a measurement is obtained [d].
Figure 1: From top to bottom, left to right, the Figure shows panels of optical, X-ray and $\gamma$-rays light curves, waveforms of the sonification as a function of time, and spectrograms of the blazar Mrk 501. The sonification is available at https://www.guijongustavo.org/datasonification/mrk501/playlist.html
(2) Mrk 1501 is a FSRQ with redshift of $z$ = 0.089338 at a distance of $\sim$ 377 Mpc (Sargent & Searle, 1970). This blazar presents short-lived $\gamma$-ray emission flares Arsioli & Polenta (2018). It is in the Pisces’s constellation, with R. A. = 00h 10m, and Dec. = +10° 58’. (Markarian et al., 1989).
For the optical, the dataset was taken from the ZTF database, from 2010 July 25 to 2021 November 25. In the case of X-rays the database comes from the Swift observatory from 2008 October 2 to 2021 March 22. For the period of 2008 August 4 to 2022 October 22, the dataset is from the Fermi database. Table 2 shows the summary of the dataset, and Figure 2 presents the light curves.
Band Records Interim Total
 [y, m, d] [d]
Optical $1255$ $11,4,0$ $4141$
X-rays $11150$ $12,5,20$ $4554$
$\gamma$-rays $5146$ $14,2,18$ $5192$
Table 2: The table shows each electromagnetic band studied for Mrk 1501: the number of records, the interim in years, months, and days [y, m, d], and the total days for which at least one measurement is taken [d].
Figure 2: From top to bottom, left to right,the Figure shows panels of optical, X-ray and $\gamma$-rays light curves, waveforms of the sonification as a function of time, and spectrograms of the blazar Mrk 1501. The sonification is available in https://www.guijongustavo.org/datasonification/mrk1501/playlist.html
(3) BL Lacerta was discovered in 1929 by Hoffmeister (1929) and Oke & Gunn (1974) measured its redshift $z$ = 0.07 which corresponds to a distance of $\sim$ 0.276 Gpc. It is in the constellation Lacertae, with R. A. = 22h 02m 43.3s, Dec. = +42° 16’ 40” (Miller et al., 1978).
The dataset in radio for this object was obtained from UMRAO from 2009 September 1 to 2012 May 21. For optical, the dataset used was from AAVSO in an interim of 1969 December 14 to 2022 March 13. For X-rays, the time interval is 2008 October 2 to 2022 November 10, taken from the Swift Observatory. And for $\gamma$-rays, the time of observation was from 2008 August 4 to 2022 March 11. Table 3 presents the summaries of the dataset and the light curves are shown in Figure 3.
Band Records Interim Total
 [y, m, d] [d]
Radio $413$ $12,5,25$ $4560$
Optical $10117$ $52,2,27$ $19082$
X-rays $100418$ $14,1,8$ $5152$
$\gamma$-rays $4892$ $13,7,7$ $4967$
Table 3: The table shows each electromagnetic band studied for BL Lacertae: the number of records, the interim in years, months, and days [y, m, d], and the total days for which at least a measurement is obtained [d].
Figure 3: From left to right, top to bottom, the Figure shows panels of radio, optical, X-ray and $\gamma$-rays light curves, waveforms of the sonification as a function of time, and spectrograms of the blazar BL Lacerta. The sonification is available in https://www.guijongustavo.org/datasonification/bllacertae/playlist.html
[…]
For AO 0235+164, the optical dataset was obtained from 1998 August 25 to 2020 September 13, and the database used was from AAVSO. For X-rays, the dataset used was from Swift from 2008 October 2 to 2017 December 7. The Fermi $\gamma$-rays database was taken from 2008 August 4 to 2020 April 5. Table 4 shows a summary of these details and Figure 4 presents the corresponding light curves.
Band Records Interim Total
 [y, m, d] [d]
Optical $196$ $22,9,17$ $8055$
X-rays $61146$ $9,2,5$ $3353$
$\gamma$-rays $4223$ $11,8,1$ $4262$
Table 4: The table shows each electromagnetic band studied for AO 0235+164: the number of records, the interim in years, months, and days [y, m, d], and the total days for which at least a measurement is obtained [d].
Figure 4: From top to bottom, left to right, the Figure shows panels of optical, X-ray and $\gamma$-rays light curves, waveforms of the sonification as a function of time, and spectrograms of the blazar AO 0235+164. The sonification is available in %****␣magallanes-mendoza.tex␣Line␣575␣****https://www.guijongustavo.org/datasonification/ao0335/playlist.html
(5) 3C 66A has a redshift of $z$ = 0.34 corresponding to $\sim$ 1.4 Gpc (Torres-Zafra et al., 2018). This BL Lac object has been observed on optical wavelengths (e. g. Böttcher et al., 2005; Lainela et al., 1999). It has R.A. = 2h 23m 12s and Dec.= 43° 0.’7 reported by Errando et al. (2009) in $\gamma$-rays.
The radio records from the Astro-geo VLBI were taken from 1997 January 11 to 2020 June 4. For optical, the dataset used was from AAVO from 1994 February 1 to 2022 February 10. The Swift X-rays database used is from 2008 October 2 to 2022 December 9. All these specifications are presented in Table 5 and the light curves are shown in Figure 5.
Band Records Interim Total
 [y, m, d] [d]
Radio $101$ $23,4,22$ $8545$
Optical $8470$ $28,0,9$ $10236$
X-rays $3163$ $14,2,7$ $5181$
$\gamma$-rays $4223$ $11,8,1$ $4262$
Table 5: The table shows each electromagnetic band studied for 3C 66A: the number of records, the interim in years, months, and days [y, m, d], and the total days for which at least a measurement is obtained [d].
Figure 5: From left to right, top to bottom, the Figure shows panels of radio, optical, X-ray and $\gamma$-rays light curves, waveforms of the sonification as a function of time, and spectrograms of the blazar 3C 66A. The sonification is available in https://www.guijongustavo.org/datasonification/3c66a/playlist.html
(6) OJ 049 has a redshift $z$ = 0.18 $\sim$, corresponding to a distance of $\sim$ 760 Mpc (Falomo, 1991). It was detected as an X-rays source (Della Ceca et al., 1990) and as a GeV $\gamma$-rays source (Fichtel, 1994; von Montigny et al., 1995; Mattox, 1998; Mukherjee et al., 1997). Its R. A. = 08h 31m 48.8s and Dec. = $04^{\circ}$ 29’ 39” (Fiorucci, M. & Tosti, G., 1996).
For this blazar the databases were AAVSO, Swift and Fermi. The corresponding time periods were from 2008 January 9 to 2020 June 23, 2002 October 9 to 2020 March 31, 2006 October 23 to 2021 December 31 and 2008 August 5 to 2020 April 6, respectively. Table 6 shows all these details and their corresponding light curves are presented in Figure 6.
Band Records Interim Total
 [y, m, d] [d]
Optical $76$ $17,5,22$ $6383$
X-rays $16111$ $15,2,8$ $5548$
$\gamma$-rays $4196$ $11,8,1$ $4262$
Table 6: The table shows each electromagnetic band studied for 0J 049: the number of records, the interim in years, months, and days [y, m, d], and the total days for which at least a measurement is obtained [d].
Figure 6: From top to bottom, left to right, the Figure shows panels of optical, X-ray and $\gamma$-rays light curves, waveforms of the sonification as a function of time, and spectrograms of the blazar OJ 049. The sonification is available in https://www.guijongustavo.org/datasonification/oj049/playlist.html
(7) PKS J2134-0153 has a redshift $z$ = 1.285, corresponding to a distance $\sim$ 5 Gpc (Truebenbach & Darling, 2017) with R. A. = 21h 34m 10.3095s and Dec. = $-01^{\circ}$ 53’ 17.238” (Fomalont et al., 2000).
For optical, the database is from ZTF with interim from 2009 October 4 to 2010 August 14. In X-rays, the dataset is from Swift from 2008 October 2 to 2014 October 8. For $\gamma$-rays, the database is Fermi, with records from 2008 August 4 to 2022 October 22. All these details are presented in Table 7. Figure 7 shows the corresponding light curves.
Band Records Interim Total
 [y, m, d] [d]
Optical $73$ $0,10,10$ $314$
X-rays $4123$ $6,0,6$ $2197$
$\gamma$-rays $4223$ $11,8,1$ $4262$
Table 7: The table shows each electromagnetic band studied for PKS J2134-0153: the number of records, the interim in years, months, and days [y, m, d], and the total days for which at least a measurement is obtained [d].
Figure 7: From top to bottom, left to right, the Figure shows panels of optical, X-ray and $\gamma$-rays light curves, waveforms of the sonification as a function of time, and spectrograms of the blazar J2134-0153. The sonification is available in https://www.guijongustavo.org/datasonification/j2134/playlist.html
(8) Mrk 421 has a R. A. = 11h 04m 27 31s and Dec. = $38^{\circ}$ 12’ 31.80”, located at a redshift $z$ = 0.031, which corresponds to a distance $\sim$ 130 Mpc) (Ulrich et al., 1975). This object shows rapid flux and polarisation variability (e. g., Fraija et al., 2017).
For optical, the dataset is from AAVSO database, from 1984 June 22 to 2020 April 13. In the case of X-rays, the database used was Swift, from 2005 March 1 to 2020 April 1. For $\gamma$-rays, the database Fermi gave the dataset from 2008 June 5 to 2020 March 8. The summary of all this is presented in Table 8. The correspond light curves are shown in Figure 8.
Band Records Interim Total
 [y, m, d] [d]
Optical $18167$ $35,9,20$ $13079$
X-rays $160346$ $15,1,0$ $5510$
$\gamma$-rays $4129$ $11,9,3$ $4294$
Table 8: The table shows each electromagnetic band studied for Mrk 421: the number of records, the interim in years, months, and days [y, m, d], and the total days for which at least a measurement is obtained [d].
Figure 8: From top to bottom, left to right, the Figure shows panels of optical, X-ray and $\gamma$-rays light curves, waveforms of the sonification as a function of time, and spectrograms of the blazar Mrk 421. The sonification is available in https://www.guijongustavo.org/datasonification/mrk421/playlist.html
(9) OJ 287 is a BL Lacerta type quasar situated at a redshift of $z$ = 0.306 $\sim$ 4 Gly $\sim$ 1 Gpc with R. A. = 08h 54m 48.87s and Dec. = +20° 06’ 30.6”. This blazar is a candidate to host a binary supermassive black hole (Valtonen et al., 2006) with a periodicity of 12 years in its optical wavelength (Shi et al., 2007).
For optical, the database was obtained from AAVSO, from 1987 September 26 to 2020 June 3. Swift provided the dataset of X-rays from 2008 November 12 to 2020 April 28. In the case of $\gamma$-rays the dataset is from Fermi, from 2008 April 4 to 2020 June 6. Table 9 show the records of this blazar. Figure 9 presents the corresponding light curves.
Band Records Interim Total
 [y, m, d] [d]
Optical $13754$ $32,8,6$ $11939$
X-rays $9464$ $11,5,16$ $4185$
$\gamma$-rays $4410$ $12,2,2$ $4446$
Table 9: The table shows each electromagnetic band studied for OJ 287: the number of records, the interim in years, months, and days [y, m, d], and the total days for which at least a measurement is obtained [d].
Figure 9: From top to bottom, left to right, the Figure shows panels of optical, X-ray and $\gamma$-rays light curves, waveforms of the sonification as a function of time, and spectrograms of the blazar OJ 287. The sonification is available in https://www.guijongustavo.org/datasonification/oj287/playlist.html
3.2 Production
Once the processed data were available, the production consisted of converting signal to sound. For this, the data was normalised to be expressed in linear units with Musical Instrument Digital Interface (MIDI) notes.
Sonification was achieved using the Parameter Mapping Sonification (PMSon) with the technique ‘one-to-one’ mapping (Grond & Berger, 2011), where the parameters are mapped to the signal domain at a tempo of 80 bpm, C3 octave, using a relative d minor scale.
All the programming was carried out using the MIDItime 1.1.3 (Mark, 2016) package using the Python 3 programming language.
3.3 Post-production
With the MIDI instrument interface of the Linux MultiMedia Studio (LMMS) software, MIDI notes were synthesised using the ZynAddSubFX plugin. This plugin is an open source synthesizer that combines different methods of audio synthesis: additive synthesis, subtractive synthesis, Fourier, and an algorithm for generating wavetable synthesis. The sound used to represent the MIDI pulses was a tinkle bell musical instrument without effects.
The PMson used was the technique ‘one-to-many’ mapping or divergent mapping, which converts the MIDI sound to tinkle bell. This method expands the scales in the ranges of the perceptual domain.
The multifrequency sonificated objects present their time series, their waveforms and their spectrograms. Their auditory result can be consulted in the following website: Datasonification Project.
3.3.1 Waveforms and spectrograms
The visual representation of the objects were represented in three ways: the light curves, the waveforms and the spectrograms. All plots are paired in the dimension of time.
For the data of each blazar, these waveforms, are a representation of a sound wave in function of time, where the amplitude represents the volume of the sound. For all studied blazars, since the datasets were normalised, then the maximum amplitude is 1.
The spectrograms provide information about the frequencies that make up a sound, in which we can appreciate the graphic representation of how the frequency varies over time. The vertical color bar to the right of the spectrogram indicates the density accumulation of points given by the signal at each point in time and frequency. In this context, density accumulation refers to how strong the signal is—either in terms of photon counts or flux—at a given energy (or frequency) and time. Warmer colors (e.g., red or yellow) represent higher densities, meaning stronger emission, while cooler colors (e.g., blue or black) indicate lower densities or weaker signals. This allows us to visualize how the energy output of the source evolves over time and across different frequencies. Understanding the frequency, as the pitch of the instrument (in this case a tinkle bell instrument), the faster the wave repeats itself (given the frequency), the higher the pitch of the note.
It is important to mention that any physical object or phenomenon that emits noise periodically (hundreds or thousands of times per second) must have a perceptible tone. However, not all sound is periodic. For example, since white noise represents static sound, it is just a uniform distribution of audible frequencies, or Gaussian distribution. Therefore, because white noise is non-periodic, it does not have a noticeable pitch.
4 Discussion
Using a bell as a sonification instrument in this part of the production enhances clarity and auditory distinction. Its resonant tone makes it easier to perceive important events in the data without blending with other sounds.
Additionally, its sharp attack (well-defined onset) allows for precise marking of specific points in time, which is particularly useful for sonifying time series or discrete events. The bell also generates harmonics that enrich the tonal representation of data variations, helping to distinguish between different values or categories.
For example, in the light curves presented here, higher flux values translate into a more intense or higher-amplitude sound that gradually fades over time, fully covering the gaps in the curve. While these changes are evident in waveforms, they may go unnoticed in light curves. However, the ear can detect abrupt shifts in intensity or volume.
This is the advantage of sonification: what escapes the eye can be perceived by the ear.
Synchronising the three plots (light curve, waveform, and spectrogram) along the time axis provides a multidimensional representation of the data. The light curve illustrates how the intensity or flux of the signal changes over time, offering a global view of the source’s luminosity or intensity. This makes it possible to identify transient events such as peaks, dips, eclipses, explosions, or periodic variations.
The waveform, on the other hand, captures the full temporal evolution of the signal with high amplitude resolution. It acts as a detailed zoom into the temporal variations that might not be visible in the light curve. This allows for the analysis of precise dynamics, such as rapid pulses, subtle oscillations, or specific shapes. It also helps to distinguish noise from real signals and correlate light curve peaks with their exact amplitude and timing.
The spectrogram decomposes the signal into its frequency components, showing how they evolve over time. It helps to identify events linked to specific frequency bands, such as peaks or oscillation modes and distinguishes noise, characterised by specific spectral patterns, from significant signals.
Combining these three plots provides a multiscale analysis:
  • The light curve offers a general view of the signal’s behaviour over time.
  • The waveform adds fine detail about amplitude variations.
  • The spectrogram reveals frequency properties, uncovering patterns not evident in the other two.
This integrated approach enables the correlation of complex events. For instance, a peak in the light curve can be examined in the waveform to determine if it represents a sudden event and in the spectrogram to check for associated frequency changes. This is relevant for identifying specific physical phenomena, such as periodic pulses, spectral bursts, or emission changes, often observed in the study of blazars or binary stars.
Moreover, this combination is key to differentiate real signals from noise. An apparent change in the light curve that turns out to be noise can be confirmed using the waveform and spectrogram if no clear patterns in amplitude or frequency are detected. For transient events or explosions, the light curve can measure the event’s total duration and peak intensity, the waveform provides details about the pulse dynamics, and the spectrogram highlights the specific frequency bands involved.
For example, in the observation of a blazar, the light curve tracks global brightness variations, the waveform reveals rapid underlying pulses invisible on a larger scale, and the spectrogram identifies harmonic components or specific spectral features.
The synchronised combination of light curve, waveform, and spectrogram provides a comprehensive and integrated view of the data: intensity, temporal dynamics, and frequency structure. This enhanced analysis, facilitates the identification of complex events and improves the interpretation of physical phenomena.
As mentioned in the Introduction, there are already several works on data sonification in astronomy. These works provide important information from the perspective of sonifying the image, since the areas with the highest light intensity can be distinguished with the ear. In addition, they have a high pedagogical value, since they allow the public to have a didactic and playful approach to astronomy.
However, these works do not show on the time line how are the variations, trends, cyclicity, periodicity, seasonality of the phenomena or of the radiative processes. In this sense, the route that we have taken is that of sonification of the time series, or light curves, and the exercise that we have carried out has been on nine active galaxies that contain a supermassive black hole in their nucleus, an accretion disk that surrounds it, and a relativistic jet, so the radiation emission from these objects is multifrequency. Thus, sonification data for the Mrk 501, Mrk 1501, Mrk 421, BL Lacerta, AO 0235+164, 3C 66A (PKS 0219+428), OJ 049 (PKS 0829+046), OJ 287, and PKS J2134-0153 blazars were presented with their associated light curves, waveforms and spectrograms.
5 Conclusions
This study presents a multifrequency exploration of blazars, combining visualisation techniques (light curves, waveforms, and spectrograms) with sonification as complementary tools for representing and analysing their multifrequency variability. The integration of these methods not only enables a richer interpretation of the data but also provides perspectives for scientific communication, particularly for visually impaired communities.
Although this approach does not aim to definitively demonstrate periodicities, the auditory representation offers an innovative way to identify potential patterns, correlate events across different time scales, and explore the complex dynamics of these astrophysical objects. This work lays the groundwork for future research to delve deeper into the potential of sonification to highlight regularities and transient phenomena in astronomical data.
This multimodal approach contributes to a more inclusive and diverse understanding of blazar behaviour, paving the way to expand scientific analysis through methods that integrate both visual and auditory perception.
The radiation processes in the studied galaxies are multifrequency, so the data we have processed for sonification correspond to radio, optical, X-ray, and gamma-ray bands. Thus, the auditor can listen to the recorded signals that are emitted due to the radiative processes that distinguish the components of the active nucleus of the galaxy. Data were obtained from seven observatories: UMRAS, AAVSO, Astrogeo-VLBI, Swift, Fermi, and ZTF.
The graphs presented for each of the blazars are specially useful for detecting changes over time, as well as getting a closer idea of sonification in visual terms. In these one can visualise what has been sounded, the changes in intensity (volume) and frequency (tone).
It is important to say that other fields in astronomy in which the sonification of data is present are in matters of education and public participation. There is a record of how sonifying images is impacting the dissemination of astronomy, and how citizen participation is increasing (Zanella et al., 2022).
The sonification of data is a way with which substantial information can be extracted, such as the search for periodicities, power increases, regularities, absences of information, etc. In addition to visualising the data, we have sonified it with the purpose of disseminating our advances to an even broader spectrum of society, at the same time that we investigate auditory events through the sense of hearing that help to distinguish periodicities.
Thus, the central interest of this work has been to disseminate the analysis of astrophysical data from a technique that is attractive to science in general, which is listening to the phenomena. In this sense, this is a work to bring the general public closer to astrophysics, from the sonification of the light curves of the active nuclei of galaxies, but also from the concern to socially include in science sectors of the population are blind or vision impaired (BVI) (Pérez-Montero, 2019).
Acknowledgements
This work was supported by a PAPIIT DGAPA-UNAM grants IN110522 and IN118325. GMG and SM acknowledge support from Secihti (378460, 26344). We thank the variable star database of observations from the AAVSO International Database contributed by observers worldwide. We thank the public data observations from the Swift data archive and the Fermi Gamma-Rays Space Telescope collaboration for the public database used in this work.
The operation of UMRAO is supported by funds from the University of Michigan Department of Astronomy, which we acknowledge too. ‘This research has made use of data from the University of Michigan Radio Astronomy Observatory which has been supported by the University of Michigan and by a series of grants from the National Science Foundation, most recently AST-0607523’.
We thank the use of archival calibrated VLBI data from the Astro-geo Center database maintained by Leonid Petrov.
This work has used observations obtained with the Samuel Oschin 48-inch Telescope at the Palomar Observatory as part of the Zwicky Transient Facility project. ZTF is supported by the National Science Foundation under Grant No. AST-1440341 and a collaboration including Caltech, IPAC, the Weizmann Institute for Science, the Oskar Klein Center at Stockholm University, the University of Maryland, the University of Washington, Deutsches Elektronen-Synchrotron and Humboldt University, Los Alamos National Laboratories, the TANGO Consortium of Taiwan, the University of Wisconsin at Mil- waukee, and Lawrence Berkeley National Laboratories. Operations are conducted by COO, IPAC, and UW. This work has made use of publicly available data from ZTF (https://irsa.ipac.caltech.edu/Missions/ztf.html).
Data Availability
The data underlying this article are available in the article and in its online supplementary material, together with the complementary webpage: https://www.guijongustavo.org/datasonification.

### v3

@@
 Consequently, graphical representation stands as a fundamental method for conducting research in the realms of science and study.
-DataViz stands as a important methodology for both scientific communication and research.
+Moreover, its ubiquity has made it not only an analytical tool but also the default and almost universal medium for communicating information across scientific, educational, and even everyday contexts, where visualisation often constitutes the first and most natural choice for conveying knowledge (Harris, 1999).
+DataViz stands as an important methodology for both scientific communication and research.
 Nonetheless, it is essential to consider a couple of factors.
 Firstly, it is important to recognise that visual communication is not accessible to individuals who are blind or have visual impairments (BVI) (Pérez-Montero, 2019).
-Secondly, relying solely on graphical representations might prove insufficient when dealing with comprehensive data analysis.
-Therefore, incorporating supplementary review methods becomes indispensable to enhance the research process.
-Taking this into account, data sonification can serve not only as a means of inclusivity in scientific communication but also as an additional tool for research (Sawe et al., 2020).
+Secondly, relying solely on graphical representations might prove insufficient when dealing with comprehensive data analysis, which involves exploring large, multidimensional datasets, detecting subtle correlations, or interpreting complex temporal and spatial patterns.
+Therefore, incorporating supplementary review methods—such as statistical modelling, sonification, and interactive data exploration—becomes indispensable to ensure a more complete and robust understanding of the data (Hearst, 2011).
+Data sonification can serve not only as a means of inclusivity in scientific communication but also as an additional tool for research (Sawe et al., 2020).
 By transforming data into sound, it offers an alternative mode of perception that complements visual representations, allowing for a deeper understanding of complex data.
-As Kramer et al.
-(1999) defines it, data sonification as “the transformation of data relations into perceived relations in an acoustic signal for the purposes of facilitating communication or interpretation” highlights sonification as a valuable supplementary instrument for comprehending information.
-Unlike the sense of vision, hearing excels in perceiving information over time, as well as recognising patterns and momentary fluctuations.
-Therefore, sonification proves to be an excellent aid in enhancing knowledge understanding (Guttman et al., 2005).
-Another definition of data sonification, provided by Hermann (2008), characterises it as a technique that takes data as its input and produces sound signals, possibly in response to optional additional stimulation or triggering.
-This broad definition encompasses any process that uses data to create auditory outputs and can be referred to as sonification, if and only if the following conditions are satisfied:
+Building on the limitations of purely visual approaches, data sonification and sound visualization have emerged as complementary methodologies for exploring and analyzing complex datasets.
+Notably, Kaper et al.
+(2000) have demonstrated how digital sound synthesis, via tools such as DIASS (Digital Instrument for Additive Sound Synthesis), combined with immersive sound visualization in virtual-reality environments like M4CAVE, can facilitate the identification of patterns, correlations, and anomalies in multidimensional scientific data.
+Their work illustrates that auditory representations not only enhance inclusivity for blind or visually impaired (BVI) individuals but also provide all researchers with an additional perceptual channel, allowing subtle temporal and structural features in the data to be perceived more effectively.
+In this context, Kramer et al.
+(1999) define data sonification as “the transformation of data relations into perceived relations in an acoustic signal for the purposes of facilitating communication or interpretation.” This definition positions sonification as a supplementary tool that can aid in understanding complex datasets.
+Research indicates that the human auditory system is particularly effective at perceiving information distributed over time and detecting temporal patterns, often outperforming vision in these tasks (McAdams & Bigand, 1993).
+Consequently, sonification proves to be an effective method for enhancing comprehension of temporal and sequential information, revealing trends, fluctuations, and correlations in datasets that might be less apparent in visual representations alone  (Guttman et al., 2005).
+Another definition of data sonification, provided by Hermann (2008), characterises it as a technique that takes data as its input and produces sound signals.
+In some cases, the generation of these sounds may be influenced by additional stimuli or triggers, which are optional and not strictly necessary for the process.
+This broad definition considers as sonification any process that uses data to create auditory output if the following conditions are satisfied:
   • The sound reflects objective properties or relations of the input data.
@@
   • The sonification is reproducible: given the same data and identical interactions (or triggers), the resulting sound has to be structurally identical.
-  • The system can intentionally be used with different data and also be used in repetition with the same data.
-Following Hermann (2008) and Vogt (2008), the “sonification is the data-dependent generation of sound, if the transformation is systematic, objective and reproducible, so that it can be used as scientific method.”
+  • The system is designed to be reusable and generalizable, meaning it can be applied to different datasets or used iteratively to verify consistency with the same dataset.
+Following Hermann (2008) and Vogt (2008), this paper adopts the definition of sonification as the data-dependent generation of sound, provided that the transformation is systematic, objective, and reproducible, so that it can be used as a scientific method.
+This definition is chosen because it emphasizes the methodological rigor required for sonification to be considered a reliable tool for scientific analysis, distinguishing it from more artistic or exploratory uses of sound.
 Numerous instances demonstrate the effectiveness of data sonification as a valuable science communication tool.
-Notable examples include its application in depicting salmon migration patterns (Hegg et al., 2018) and visualising fluctuations in brainwave activity (Parvizi et al., 2018).
-The Geiger counter represents yet another compelling case in point.
-In the realm of astronomy, early instances of data sonification trace to Jansky (1933), when detected radio waves emanating from the centre of the Milky Way were analysed by telephone communication noise.
-Three decades later, Penzias & Wilson (1965) identified the interference in their observations as originating from cosmic microwave background radiation, marking a significant discovery in the field (Kellermann et al., 2020).
+For the purposes of this paper, science communication refers to the process of conveying scientific information and insights to diverse audiences—researchers, students, and the general public—in a manner that is both understandable and engaging.
+Notable examples include its application in depicting salmon migration patterns  (Hegg et al., 2018) and visualising fluctuations in brainwave activity (Parvizi et al., 2018).Another compelling case is the Geiger counter, a device used to detect and measure ionizing radiation.
+In traditional use, the counter produces audible clicks corresponding to radiation events, effectively sonifying the underlying radioactive activity.
+This auditory representation allows listeners to perceive variations in radiation intensity over time, providing both an immediate, intuitive understanding of the data and a practical tool for real-time monitoring.
+In the realm of astronomy, early connections between sound and scientific data can be traced to Jansky (1933), who first detected radio waves emanating from the centre of the Milky Way and identified them through the auditory noise in telephone communications.
+While Penzias & Wilson (1965) three decades later detected the cosmic microwave background radiation using radio instrumentation, their work exemplifies how radio data—initially experienced as noise—would later inspire the development of auditory and sonification approaches in astronomy (Kellermann et al., 2020).
 Another example is that of Morgan et al.
 (1997) who took the data of the X-ray emission of the black hole GRS 1915+105 and translated it into audio signals, allowing us to “hear” the accretion disk from it Masetti (2013).
-Also Chandra (2003) has sonified data and released “Hears” of a black hole for the first time.
+Also Chandra (2003) has sonified data and released “Chandra ‘Hears’ of a black hole for the first time”.
 Abbott et al.
 (2016) documented the simultaneous detection of a brief signal by two gravitational wave detectors, which resulted from the fusion of two black holes.
-The researchers converted the data from the instant of the collision, which they dubbed a “chirp.” This term was chosen because certain events producing gravitational waves exhibit auditory similarities to the chirping sounds of birds (LIGO Collaboration, 2016, see https://youtu.be/QyDcTbR-kEA).
-Furthermore, Berti (2016) published an article entitled The First Sounds of Merging Black Holes, shedding light on the initial auditory cues of black hole mergers.
-In recent times, the National Aeronautics and Space Administration (NASA) has unveiled sonification imagery derived from optical data captured by the Hubble Space Telescope (NASA, 2022) and X-ray data observed by the Chandra X-ray Observatory (CXO) (Chandra, 2020).
+The researchers converted the data from the instant of the collision, which they dubbed a “chirp,” a term chosen because the sonified signal exhibits auditory similarities to the chirping sounds of birds (LIGO Collaboration, 2016, see https://youtu.be/QyDcTbR-kEA), even though this sound arises solely from the sonification method and not from any actual audible event.
+Furthermore, Berti (2016) published an article entitled “The First Sounds of Merging Black Holes”, shedding light on the initial auditory cues of black hole mergers.
+In recent times, the National Aeronautics and Space Administration (NASA) has unveiled sonification of imagery derived from optical data captured by the Hubble Space Telescope (NASA, 2022) and X-ray data observed by the Chandra X-ray Observatory (CXO) (Chandra, 2020).
+Recent advances in astronomical sonification have demonstrated the potential of combining auditory and visual data representations.
+For instance, Tucker Brown et al.
+(2022a) assessed Astronify (Brasseur et al., 2023), a prototype tool integrated within the MAST archive that enables sonification of evenly sampled light curves.
+Their study evaluated the effectiveness of sonification for signal detection at various signal-to-noise ratios, showing that auditory representations can aid in identifying high-SNR transits and complement traditional visual inspection.
+More recently, Huppenkothen et al.
+(2023) presented the Sonified Hertzsprung–Russell Diagram, which transforms stellar time series into sound, encoding physically meaningful features that preserve astrophysical differences between stars through both visual and auditory media.
+Our work extends these approaches by applying sonification to multiwavelength light curves of blazars, integrating light curves, waveforms, and spectrograms to explore temporal evolution across distinct energy regimes (optical, X-ray, and $\Gamma$-ray).
+This multifrequency perspective highlights variability patterns in active galactic nuclei and demonstrates how sonification can support both scientific interpretation and public engagement in high-energy astrophysics.
 In the context of communicating science to the BVI community, especially in the realm of astrophysics and black holes (BHs) this work sonifies data from nine blazars: Mrk 501, Mrk 1501, Mrk 421, BL Lacerta, AO 0235+164, 3C 66A (PKS 0219+428), OJ 049 (PKS 0829+046), OJ 287, and PKS J2134-0153.
-[…]
-With this background, in Section 2 we show the way to obtain the multifrequency data (radio, optical, X-ray, and $\gamma$-ray).
+These objects have been reported in the literature as exhibiting quasi-periodic variability in different electromagnetic bands.
+By sonifying their light curves, our aim is to explore whether auditory perception can aid in identifying, characterising, and communicating such periodicities and variability patterns.
+Further details are given in Appendices A and B.
+The selection of blazars analyzed in this study was based on three main criteria: (i) the availability of long-term, sampled light curves across multiple wavelengths (radio, optical, X-ray, and gamma-ray), (ii) documented evidence of significant variability on different timescales, and (iii) their inclusion in public monitoring programs that ensure data accessibility and reproducibility.
+The chosen sources—such as Mrk 421, Mrk 501, and 3C 273—are among the most extensively observed and best-characterized blazars in the literature, known for their pronounced flux variations and strong multi-band correlations (Abdo et al., 2011).
+These properties make them ideal test cases for exploring how sonification can enhance the perception and interpretation of temporal patterns in multifrequency datasets.
+The ensemble of data employed in this study was selected to encompass a broad range of energy bands (radio, optical, X-ray, and $\gamma$-ray) in order to capture the multifrequency variability characteristic of blazars.
+While the temporal coverage is not fully simultaneous across all bands, the datasets from the Swift and Fermi observatories are largely contemporaneous, as both missions were launched in 2008, providing overlapping X-ray and gamma-ray observations over extended periods.
+This temporal alignment allows for meaningful cross-band comparisons in several sources.
+For other frequency ranges, such as optical and radio, the data were chosen to complement these high-energy observations, offering insights into long-term variability trends.
+From a sonification perspective, these differences in time sampling and frequency response enrich the auditory experience, allowing users to perceive asynchronous variability patterns, phase shifts, and multiwavelength correlations through sound.
+This sonification serves a dual goal: it enables BVI audiences to directly engage with astrophysical data, perceiving temporal patterns and variability that are typically conveyed visually, while also providing astronomers with an analytical tool to detect trends and anomalies in light curves through auditory perception, thereby complementing conventional visual analysis methods.
+Although this study does not aim to definitively demonstrate periodicities, the auditory representation provides a means to detect potential patterns, correlate events across different time scales, and explore the complex dynamics of these astrophysical objects.
+By demonstrating how sonification can complement traditional visualizations, this work establishes a methodological foundation for future research, including systematic searches for periodicities, transient events, and multiwavelength correlations in astronomical data.
+This multimodal approach enables a more inclusive and perceptually rich exploration of the variability of blazars for both sighted and visually impaired audiences.
+Blazars are extragalactic sources of intense luminosity, with energy outputs ranging from $10^{41}$ to $10^{47}$ ergs s-1 (Blandford et al., 2019), whose relativistic jets are oriented close to the observer’s line of sight (Ulrich et al., 1997).
+This geometric alignment results in strong Doppler boosting, which amplifies the observed emission and produces rapid variability across the electromagnetic (EM) spectrum.
+Their radiation is predominantly non-thermal and arises from two main mechanisms: synchrotron emission, which dominates the radio-to-optical range, and inverse Compton scattering, responsible for X-ray and gamma-ray photons (Padovani et al., 2012).Their pronounced variability over a wide range of timescales makes them particularly suitable for this study, where they serve as test cases to assess the potential of sonification both as an analytical tool and as a medium for science communication.
+The variability observed at different wavelengths reflects physical changes in distinct regions of the jet — for instance, fluctuations in magnetic field strength, particle acceleration, or shock propagation — leading to complex temporal structures that can range from minutes to years.
+Because these variations encode valuable information about the underlying astrophysical processes, blazars represent ideal candidates for data sonification.
+Transforming their multi-band light curves into sound provides an alternative perceptual channel for detecting periodicities, correlations, or transient events that might be less evident through visual inspection alone.
+With this background, in Section 2 we describe how we obtained the multifrequency data (radio, optical, X-ray, and $\gamma$-ray).
 Section 3 shows the sonification procedure: pre-production, production, and post-production.
@@
 2.1 Radio
-The radio datasets for this work were obtained from two data bases: the University of Michigan Radio Astronomy Observatory (UMRAO) [^1] https://dept.astro.lsa.umich.edu/datasets/umrao.php, and the Astrogeo Very-long-baseline interferometry (Astrogeo-VLBI).
-The 26-meter paraboloid UMRAO monitoring in the bands: 4.8, 8, and 14.5 GHz (Aller et al., 1985).
-It is focused on the study of total flux density and linear polarisation of active extragalactic objects [^2] https://lsa.umich.edu/astro/facilities/data-sets.html.
-Radio Fundamental Catalog (RFC), of the Astro-geo Center, maintained by L.
-Petrov333 http://astrogeo.org/rfc provides precise positions with milli-arcseconds accuracies, maps, and estimates of correlated flux densities at baselines 1000–8000 km of thousands of compact radio sources.
+The radio datasets for this work were obtained from two data bases: the University of Michigan Radio Astronomy Observatory (UMRAO) [^1] https://dept.astro.lsa.umich.edu/datasets/umrao.php, and the Astrogeo Very-long-baseline interferometry (Astrogeo-VLBI) database.
+From UMRAO, we obtained long-term monitoring light curves of total flux density and linear polarization at 4.8, 8.0, and 14.5 GHz measured with the 26 m radio telescope (Aller et al., 1985).
+For the sonification, data points from the three frequencies were not averaged or combined interchangeably; instead, they were arranged in temporal order while preserving their associated frequency tags.
+This approach allowed us to create a continuous auditory sequence reflecting the multifrequency monitoring cadence, where each tone corresponds to an observation at a specific band.
+The selected frequencies are those used in UMRAO’s long-term variability program, optimized for compact radio sources.
+This structure made it possible to represent the variability across frequencies and identify temporal patterns or irregular sampling intervals through sound.
+[^2] https://lsa.umich.edu/astro/facilities/data-sets.html The Astrogeo-VLBI data, corresponding to a frequency of 8.4 GHz (Petrov & Kovalev, 2025), provides complementary spatial information through very long baseline interferometry.
+The Radio Fundamental Catalog (RFC), of the Astro-geo Center, maintained by L.
+Petrov [^3] http://astrogeo.org/rfc provides precise positions with milli-arcseconds accuracies, maps, and estimates of correlated flux densities for thousands of compact radio sources.
+The mentioned baselines of 1000–8000 km mentioned in the catalog (RFC) correspond to the distances between the radio telescopes in the very long baseline interferometry (VLBI) network.
+Longer baselines provide higher angular resolution, allowing for finer spatial details of the sources.
+These datasets also include light curves provided in tabular format (flux density vs.
+time), which are suitable for variability analysis.
 2.2 Optical
 For optical, the databases used were: American Association of Variable Star Observers (AAVSO) [^4] https://www.aavso.org/ and Zwicky Transient Facility (ZTF) [^5] https://www.ztf.caltech.edu/.
+From AAVSO, we obtained publicly available photometric light curves of blazars in different optical filters, contributed by both amateur and professional astronomers.
+The data are distributed as tab-delimited American Standard Code for Information Interchange (ASCII) tables containing Julian Date, magnitude, filter, and error.
 The optical AAVSO is a public database founded in 1910 and offers observations of variable stars collected and archived for world-wide access in collaboration with amateur and professional astronomers.
@@
 2.3 X-rays
-For the X-rays light curve, we use the Swift database with energies in the range $0.3-10\,\text{keV}$.
-It contains the X-Ray Telescope (XRT) with two important characteristics that makes it important for observations: low background and a constant point spread function across the field of view (Moretti et al., 2005).
+For the X-ray light curves, we used data from the Swift database covering energies in the range 0.3–10 keV.
+The X-Ray Telescope (XRT) has two key characteristics that are crucial for this analysis: a low background and a nearly constant point spread function (PSF) across the field of view.
+A constant PSF ensures that the spatial distribution of X-ray photons from the source is measured uniformly across the detector, reducing systematic effects when constructing light curves and enabling more reliable detection of flux variations (Moretti et al., 2005).
+The selected blazars are relatively X-ray bright, providing sufficient photon counts to generate meaningful temporal variations for both visualization and sonification.
+The Swift database provides background-subtracted light curves in Flexible Image Transport System (FITS) and ASCII table formats, including time, count rate, and errors.
+These products were directly used in our analysis with additional reduction.
 2.4 Gamma-rays
-The public database of $\gamma$-rays used was of The Fermi Large Area Telescope (LAT).
+For the gamma-ray regime, we used data from the Fermi Large Area Telescope (LAT) in the 100 MeV–300 GeV range.
+Specifically, we used binned light curves in FITS files from the public Fermi-LAT Light Curve Repository, which provides fluxes integrated over pre-defined time bins (typically weekly) in tabular format (flux vs.
+time with associated errors).
+These products are already processed through the LAT standard analysis pipeline FERMITOOLS [^6] https://fermi.gsfc.nasa.gov/ssc/data/analysis/software/ and suitable for variability studies.
 This instrument observes the sky for $\sim$ 3 hours and covers $\gtrsim$ 20% of it for each observation.
-It has a crystal calorimeter for energy measurement, an anti-coincidence detector to distinguish the background of charged particles and high-resolution converter tracker for direction measurement of the incident gamma rays (Atwood et al., 2009).
-The $\gamma$-rays fluxes collected are in the range $100\,\text{MeV}-300\text{GeV}$.
+It has a crystal calorimeter for energy measurement, an anti-coincidence detector to distinguish the background of charged particles and high-resolution converter tracker for direction measurement of the incident gamma-rays (Atwood et al., 2009).
 3 Sonification
-Sonifying the data for each of the blazar frequencies was carried out in three stages: pre-production, production and post-production.
+Sonifying the data for each of the blazar frequencies was carried out in three stages: pre-production, production and post-production detailed in the next section.
 3.1 Pre-production
-Sonification pre-production basically focuses on data pre-processing.
-In the case of the radio and optical bands, the data was processed in such a way that any anomaly, disturbance or irrelevance of the data could be detected, eliminated, corrected or transformed.
-For this task the AWK pattern scanning and processing language was used.
-In the case of X-rays, it was necessary to reduce the data and for this, the HEASoft [^6] https://heasarc.gsfc.nasa.gov/docs/software/lheasoft/ software was used.
-The data processing corresponding to the photons, the position and the orientation of the satellite is available on Flexible Image Transport System format files (FITS).
-For the pre-production of sonification in the $\gamma$-rays, the data reduction was possible with the Fermitools [^7] https://fermi.gsfc.nasa.gov/ssc/data/analysis/software/ software.
-For this, it was necessary to know the main relevant parameters of the blazars such as: R.A., Dec., Region of Interest (ROI) and start and end time of the data processing (see e.g Magallanes-Guijón, 2020).
-The process of this reduction followed the procedure detailed by Cabrera et al.
-(2013).
 […]
-3.1.1 Light curves
-Each blazar selected had the following specifications:
-(1) Mrk 501, has a redshift $z$ = 0.034 that corresponds to a distance $\sim$ 140 Mpc and R.
-A.
-= 16h 53m 52.2s, Dec.
-= +39° 45’ 37”.
-It has been monitored in radio (Richards et al., 2011), optical (Smith et al., 2009), X-rays (Abdo et al., 2011), and $\gamma$-rays (Dorner et al., 2017).
-This object has several periodicities reported (Bhatta, 2019).
-Recently Magallanes-Guijón & Mendoza (2024) have reported a periodicity $\sim$ 224 days.
-The light curve in optical was built using the database from 1998 June 24 to 2021 September 12.
-The Swift database is from 2008 October 2 to 2021 October 8.
-And the $\gamma$-rays, covers a time interval from 2008 August 04 up to 2020 March 13.
-Table 1 shows specific information about the records and the light curve is presented in Figure 1.
-Band Records Interim Total
-  [y, m, d] [d]
-Optical $11,849$ $23,2,17$ $8441$
-X-rays $28,000$ $12,7,9$ $4607$
-$\gamma$-rays $4,199$ $11,7,9$ $4239$
-Table 1: The table shows each electromagnetic band studied for Mrk 501: the number of records, the interim in years, months, and days [y, m, d], and the total days for which at least a measurement is obtained [d].
+  • Eliminated if the value was clearly spurious (e.g., magnitude far outside the instrument’s dynamic range, corrupted entries).
+  • Corrected if a small inconsistency could be addressed (e.g., date formatting or duplicated entries consolidated).
+  • Transformed if a re-scaling or homogenization was required (e.g., converting magnitudes to flux densities, or aligning different photometric systems).
+This preprocessing was performed using the AWK pattern scanning and processing language, which allows filtering, restructuring, and formatting the data into consistent light curve tables.
+For the X-ray data, a full reduction was necessary.
+Raw event files were processed with the HEASoft [^7] https://heasarc.gsfc.nasa.gov/docs/software/lheasoft/ package, which applies the standard calibration pipelines (e.g., bias subtraction, exposure corrections, effective area, and deadtime corrections).
+The photon event information, together with spacecraft attitude (position and orientation), is stored in FITS files, which serve as the standard format for astronomical data exchange and allow reproducibility of the analysis.
+For the $\gamma$-ray data, reduction was performed using the Fermitools software suite.
+The procedure required specifying the relevant astrophysical parameters of each blazar—Right Ascension (R.A.), Declination (Dec.), Region of Interest (ROI), and the start and end times of the data processing—following the standard pipeline described by Cabrera et al.
+(2013) and Magallanes-Guijón (2020).
+[…]
 Figure 1: From top to bottom, left to right, the Figure shows panels of optical, X-ray and $\gamma$-rays light curves, waveforms of the sonification as a function of time, and spectrograms of the blazar Mrk 501.
 The sonification is available at https://www.guijongustavo.org/datasonification/mrk501/playlist.html
-(2) Mrk 1501 is a FSRQ with redshift of $z$ = 0.089338 at a distance of $\sim$ 377 Mpc (Sargent & Searle, 1970).
-This blazar presents short-lived $\gamma$-ray emission flares Arsioli & Polenta (2018).
-It is in the Pisces’s constellation, with R.
-A.
-= 00h 10m, and Dec.
-= +10° 58’.
-(Markarian et al., 1989).
-For the optical, the dataset was taken from the ZTF database, from 2010 July 25 to 2021 November 25.
-In the case of X-rays the database comes from the Swift observatory from 2008 October 2 to 2021 March 22.
-For the period of 2008 August 4 to 2022 October 22, the dataset is from the Fermi database.
-Table 2 shows the summary of the dataset, and Figure 2 presents the light curves.
-Band Records Interim Total
- [y, m, d] [d]
-Optical $1255$ $11,4,0$ $4141$
-X-rays $11150$ $12,5,20$ $4554$
-$\gamma$-rays $5146$ $14,2,18$ $5192$
-Table 2: The table shows each electromagnetic band studied for Mrk 1501: the number of records, the interim in years, months, and days [y, m, d], and the total days for which at least one measurement is taken [d].
 Figure 2: From top to bottom, left to right,the Figure shows panels of optical, X-ray and $\gamma$-rays light curves, waveforms of the sonification as a function of time, and spectrograms of the blazar Mrk 1501.
 The sonification is available in https://www.guijongustavo.org/datasonification/mrk1501/playlist.html
-(3) BL Lacerta was discovered in 1929 by Hoffmeister (1929) and Oke & Gunn (1974) measured its redshift $z$ = 0.07 which corresponds to a distance of $\sim$ 0.276 Gpc.
-It is in the constellation Lacertae, with R.
-A.
-= 22h 02m 43.3s, Dec.
-= +42° 16’ 40” (Miller et al., 1978).
-The dataset in radio for this object was obtained from UMRAO from 2009 September 1 to 2012 May 21.
-For optical, the dataset used was from AAVSO in an interim of 1969 December 14 to 2022 March 13.
-For X-rays, the time interval is 2008 October 2 to 2022 November 10, taken from the Swift Observatory.
-And for $\gamma$-rays, the time of observation was from 2008 August 4 to 2022 March 11.
-Table 3 presents the summaries of the dataset and the light curves are shown in Figure 3.
-Band Records Interim Total
- [y, m, d] [d]
-Radio $413$ $12,5,25$ $4560$
-Optical $10117$ $52,2,27$ $19082$
-X-rays $100418$ $14,1,8$ $5152$
-$\gamma$-rays $4892$ $13,7,7$ $4967$
-Table 3: The table shows each electromagnetic band studied for BL Lacertae: the number of records, the interim in years, months, and days [y, m, d], and the total days for which at least a measurement is obtained [d].
 Figure 3: From left to right, top to bottom, the Figure shows panels of radio, optical, X-ray and $\gamma$-rays light curves, waveforms of the sonification as a function of time, and spectrograms of the blazar BL Lacerta.
 The sonification is available in https://www.guijongustavo.org/datasonification/bllacertae/playlist.html
-[…]
-For AO 0235+164, the optical dataset was obtained from 1998 August 25 to 2020 September 13, and the database used was from AAVSO.
-For X-rays, the dataset used was from Swift from 2008 October 2 to 2017 December 7.
-The Fermi $\gamma$-rays database was taken from 2008 August 4 to 2020 April 5.
-Table 4 shows a summary of these details and Figure 4 presents the corresponding light curves.
-Band Records Interim Total
- [y, m, d] [d]
-Optical $196$ $22,9,17$ $8055$
-X-rays $61146$ $9,2,5$ $3353$
-$\gamma$-rays $4223$ $11,8,1$ $4262$
-Table 4: The table shows each electromagnetic band studied for AO 0235+164: the number of records, the interim in years, months, and days [y, m, d], and the total days for which at least a measurement is obtained [d].
 Figure 4: From top to bottom, left to right, the Figure shows panels of optical, X-ray and $\gamma$-rays light curves, waveforms of the sonification as a function of time, and spectrograms of the blazar AO 0235+164.
-The sonification is available in %****␣magallanes-mendoza.tex␣Line␣575␣****https://www.guijongustavo.org/datasonification/ao0335/playlist.html
-(5) 3C 66A has a redshift of $z$ = 0.34 corresponding to $\sim$ 1.4 Gpc (Torres-Zafra et al., 2018).
-This BL Lac object has been observed on optical wavelengths (e.
-g.
-Böttcher et al., 2005; Lainela et al., 1999).
-It has R.A.
-= 2h 23m 12s and Dec.= 43° 0.’7 reported by Errando et al.
-(2009) in $\gamma$-rays.
-The radio records from the Astro-geo VLBI were taken from 1997 January 11 to 2020 June 4.
-For optical, the dataset used was from AAVO from 1994 February 1 to 2022 February 10.
-The Swift X-rays database used is from 2008 October 2 to 2022 December 9.
-All these specifications are presented in Table 5 and the light curves are shown in Figure 5.
-Band Records Interim Total
- [y, m, d] [d]
-Radio $101$ $23,4,22$ $8545$
-Optical $8470$ $28,0,9$ $10236$
-X-rays $3163$ $14,2,7$ $5181$
-$\gamma$-rays $4223$ $11,8,1$ $4262$
-Table 5: The table shows each electromagnetic band studied for 3C 66A: the number of records, the interim in years, months, and days [y, m, d], and the total days for which at least a measurement is obtained [d].
+The sonification is available in https://www.guijongustavo.org/datasonification/ao0335/playlist.html
 Figure 5: From left to right, top to bottom, the Figure shows panels of radio, optical, X-ray and $\gamma$-rays light curves, waveforms of the sonification as a function of time, and spectrograms of the blazar 3C 66A.
 The sonification is available in https://www.guijongustavo.org/datasonification/3c66a/playlist.html
-(6) OJ 049 has a redshift $z$ = 0.18 $\sim$, corresponding to a distance of $\sim$ 760 Mpc (Falomo, 1991).
-It was detected as an X-rays source (Della Ceca et al., 1990) and as a GeV $\gamma$-rays source (Fichtel, 1994; von Montigny et al., 1995; Mattox, 1998; Mukherjee et al., 1997).
-Its R.
-A.
-= 08h 31m 48.8s and Dec.
-= $04^{\circ}$ 29’ 39” (Fiorucci, M.
-& Tosti, G., 1996).
-For this blazar the databases were AAVSO, Swift and Fermi.
-The corresponding time periods were from 2008 January 9 to 2020 June 23, 2002 October 9 to 2020 March 31, 2006 October 23 to 2021 December 31 and 2008 August 5 to 2020 April 6, respectively.
-Table 6 shows all these details and their corresponding light curves are presented in Figure 6.
-Band Records Interim Total
- [y, m, d] [d]
-Optical $76$ $17,5,22$ $6383$
-X-rays $16111$ $15,2,8$ $5548$
-$\gamma$-rays $4196$ $11,8,1$ $4262$
-Table 6: The table shows each electromagnetic band studied for 0J 049: the number of records, the interim in years, months, and days [y, m, d], and the total days for which at least a measurement is obtained [d].
 Figure 6: From top to bottom, left to right, the Figure shows panels of optical, X-ray and $\gamma$-rays light curves, waveforms of the sonification as a function of time, and spectrograms of the blazar OJ 049.
 The sonification is available in https://www.guijongustavo.org/datasonification/oj049/playlist.html
-(7) PKS J2134-0153 has a redshift $z$ = 1.285, corresponding to a distance $\sim$ 5 Gpc (Truebenbach & Darling, 2017) with R.
-A.
-= 21h 34m 10.3095s and Dec.
-= $-01^{\circ}$ 53’ 17.238” (Fomalont et al., 2000).
-For optical, the database is from ZTF with interim from 2009 October 4 to 2010 August 14.
-In X-rays, the dataset is from Swift from 2008 October 2 to 2014 October 8.
-For $\gamma$-rays, the database is Fermi, with records from 2008 August 4 to 2022 October 22.
-All these details are presented in Table 7.
-Figure 7 shows the corresponding light curves.
-Band Records Interim Total
- [y, m, d] [d]
-Optical $73$ $0,10,10$ $314$
-X-rays $4123$ $6,0,6$ $2197$
-$\gamma$-rays $4223$ $11,8,1$ $4262$
-Table 7: The table shows each electromagnetic band studied for PKS J2134-0153: the number of records, the interim in years, months, and days [y, m, d], and the total days for which at least a measurement is obtained [d].
 Figure 7: From top to bottom, left to right, the Figure shows panels of optical, X-ray and $\gamma$-rays light curves, waveforms of the sonification as a function of time, and spectrograms of the blazar J2134-0153.
 The sonification is available in https://www.guijongustavo.org/datasonification/j2134/playlist.html
-(8) Mrk 421 has a R.
-A.
-= 11h 04m 27 31s and Dec.
-= $38^{\circ}$ 12’ 31.80”, located at a redshift $z$ = 0.031, which corresponds to a distance $\sim$ 130 Mpc) (Ulrich et al., 1975).
-This object shows rapid flux and polarisation variability (e.
-g., Fraija et al., 2017).
-For optical, the dataset is from AAVSO database, from 1984 June 22 to 2020 April 13.
-In the case of X-rays, the database used was Swift, from 2005 March 1 to 2020 April 1.
-For $\gamma$-rays, the database Fermi gave the dataset from 2008 June 5 to 2020 March 8.
-The summary of all this is presented in Table 8.
-The correspond light curves are shown in Figure 8.
-Band Records Interim Total
- [y, m, d] [d]
-Optical $18167$ $35,9,20$ $13079$
-X-rays $160346$ $15,1,0$ $5510$
-$\gamma$-rays $4129$ $11,9,3$ $4294$
-Table 8: The table shows each electromagnetic band studied for Mrk 421: the number of records, the interim in years, months, and days [y, m, d], and the total days for which at least a measurement is obtained [d].
 Figure 8: From top to bottom, left to right, the Figure shows panels of optical, X-ray and $\gamma$-rays light curves, waveforms of the sonification as a function of time, and spectrograms of the blazar Mrk 421.
 The sonification is available in https://www.guijongustavo.org/datasonification/mrk421/playlist.html
-(9) OJ 287 is a BL Lacerta type quasar situated at a redshift of $z$ = 0.306 $\sim$ 4 Gly $\sim$ 1 Gpc with R.
-A.
-= 08h 54m 48.87s and Dec.
-= +20° 06’ 30.6”.
-This blazar is a candidate to host a binary supermassive black hole (Valtonen et al., 2006) with a periodicity of 12 years in its optical wavelength (Shi et al., 2007).
-For optical, the database was obtained from AAVSO, from 1987 September 26 to 2020 June 3.
-Swift provided the dataset of X-rays from 2008 November 12 to 2020 April 28.
-In the case of $\gamma$-rays the dataset is from Fermi, from 2008 April 4 to 2020 June 6.
-Table 9 show the records of this blazar.
-Figure 9 presents the corresponding light curves.
-Band Records Interim Total
- [y, m, d] [d]
-Optical $13754$ $32,8,6$ $11939$
-X-rays $9464$ $11,5,16$ $4185$
-$\gamma$-rays $4410$ $12,2,2$ $4446$
-Table 9: The table shows each electromagnetic band studied for OJ 287: the number of records, the interim in years, months, and days [y, m, d], and the total days for which at least a measurement is obtained [d].
 Figure 9: From top to bottom, left to right, the Figure shows panels of optical, X-ray and $\gamma$-rays light curves, waveforms of the sonification as a function of time, and spectrograms of the blazar OJ 287.
@@
 Once the processed data were available, the production consisted of converting signal to sound.
-For this, the data was normalised to be expressed in linear units with Musical Instrument Digital Interface (MIDI) notes.
-Sonification was achieved using the Parameter Mapping Sonification (PMSon) with the technique ‘one-to-one’ mapping (Grond & Berger, 2011), where the parameters are mapped to the signal domain at a tempo of 80 bpm, C3 octave, using a relative d minor scale.
-All the programming was carried out using the MIDItime 1.1.3 (Mark, 2016) package using the Python 3 programming language.
+For this purpose, each dataset was first normalized using a linear min-max scaling:
+$\textrm{normalized value}=\frac{\text{flux}-\min(\text{flux})}{\max(\text{flux})-\min(\text{flux})},$ (1)
+where flux is the measured brightness at a given time, and $\min(\text{flux})$ and $\max(\text{flux})$ correspond to the minimum and maximum flux values in the dataset.
+This procedure ensures that all data points are rescaled to the $[0,1]$ interval, facilitating their subsequent mapping into musical notes.
+The normalized values were then mapped to Musical Instrument Digital Interface (MIDI) notes using Parameter Mapping Sonification (PMSon) with the ‘one-to-many’ mapping technique (Grond & Berger, 2011).
+In this approach, each normalized data point triggers a primary note (encoding flux as pitch) and a set of characteristic overtones, producing a richer auditory representation.
+Additionally, the amplitude (loudness) of each note is directly tied to the flux, meaning that the same physical quantity is represented both in pitch and amplitude.
+This dual encoding enhances the perception of variations, trends, and transient events in the light curves.
+To map the normalized values to the notes of the scale, the interval [0, 1] was divided into seven equal bins, each corresponding uniquely to one of the seven notes of the D minor scale within a single octave.
+For example, a normalized value of 0.0 corresponds to the first note (D), a value of 0.5 corresponds to the fourth note (G), and a value of 1.0 to the highest note (C) in the octave.
+The sonification was implemented in the C3 octave, with a tempo of 80 beats per minute (bpm), where each note represents one time step of the light curve.
+In this context, the 80 bpm provides a temporal spacing between successive notes, corresponding to the time evolution of the observed flux.
+Different octave ranges were assigned to each wavelength band during post-production.
+This choice does not imply any physical relationship between electromagnetic wavelength and pitch; rather, it was a perceptual design decision intended to enhance audibility and facilitate differentiation between bands in the multiwavelength sonification.
+Because the data span different time intervals and have varying sampling rates across wavelength bands, the light curves were rescaled in time for the sonification.
+Successive data points were mapped to fixed note durations at 80 bpm, representing the relative temporal evolution of each dataset.
+Absolute timing (MJD) is not preserved in the sonification; instead, temporal structures and patterns within each wavelength are maintained.
+This allows auditory perception of variability in each band, while the differences in time span and sampling between bands provide complementary insights.
+Key intervals, including known quasi-periodicities, are annotated in the figures and on the project web page to guide the listener in interpreting the sonified data.
+The programming was carried out using the MIDItime 1.1.3 package (Mark, 2016) in Python 3.
+Data uncertainties were not directly encoded in the sonification, although they informed the preprocessing stage by guiding the removal or correction of anomalous points prior to mapping.
 3.3 Post-production
@@
 The sound used to represent the MIDI pulses was a tinkle bell musical instrument without effects.
-The PMson used was the technique ‘one-to-many’ mapping or divergent mapping, which converts the MIDI sound to tinkle bell. This method expands the scales in the ranges of the perceptual domain.
-The multifrequency sonificated objects present their time series, their waveforms and their spectrograms. Their auditory result can be consulted in the following website: Datasonification Project.
+The tinkle bell produces a percussive, bell-like sound with a fundamental pitch and a set of inharmonic overtones (i.e., partials that are not exact integer multiples of the fundamental frequency).
+Each note has a rapid attack and an exponential decay determined by the instrument’s envelope.
+The spectral components scale proportionally with the input pitch, so higher notes produce perceptually brighter sounds.
+This instrument was chosen because its clear and transient timbre and distinctive spectral structure allow individual data points to be audibly distinguished, even in dense sequences.
+Parameter Mapping Sonification (PMSon) was applied using the ‘one-to-many’ mapping technique.
+In this approach, a single normalized data point generates multiple simultaneous partials based on the tinkle bell’s harmonic structure.
+Specifically, each normalized data value is first mapped to a primary MIDI note; this note then triggers the instrument’s characteristic overtones, producing a richer auditory representation and expanding the perceptual pitch content.
+To ensure perceptual separation between wavelength regimes (radio, optical, X-ray, $\gamma$-ray), each band was assigned to a distinct octave range via pitch offsets applied after the mapping stage.
+This strategy preserves the relative variations within each light curve while allowing the different regimes to occupy clearly distinguishable pitch regions.
+Each note has a constant duration of one beat at 80 bpm, corresponding to the temporal spacing defined in the sonification.
+The tinkle bell instrument was configured with an exponential decay of approximately 0.5 seconds, allowing individual notes to be clearly perceived while avoiding excessive overlap.
+Temporal overlap naturally occurs if data points are closer than the beat interval.
+Gaps in the light curves are represented by silent intervals of corresponding length, preserving the temporal structure of the original data.
+The multifrequency sonified objects are illustrated in figures showing:
+  1. 1.
+    the original light curve (time series),
+  2. 2.
+    the sonified waveform, and
+  3. 3.
+    the corresponding spectrogram, allowing visualization of both temporal and frequency content.
+The auditory results of these sonifications can be accessed at the Datasonification Project website.
 3.3.1 Waveforms and spectrograms
-The visual representation of the objects were represented in three ways: the light curves, the waveforms and the spectrograms. All plots are paired in the dimension of time.
-For the data of each blazar, these waveforms, are a representation of a sound wave in function of time, where the amplitude represents the volume of the sound.
-For all studied blazars, since the datasets were normalised, then the maximum amplitude is 1.
-The spectrograms provide information about the frequencies that make up a sound, in which we can appreciate the graphic representation of how the frequency varies over time.
-The vertical color bar to the right of the spectrogram indicates the density accumulation of points given by the signal at each point in time and frequency.
-In this context, density accumulation refers to how strong the signal is—either in terms of photon counts or flux—at a given energy (or frequency) and time.
-Warmer colors (e.g., red or yellow) represent higher densities, meaning stronger emission, while cooler colors (e.g., blue or black) indicate lower densities or weaker signals.
-This allows us to visualize how the energy output of the source evolves over time and across different frequencies.
-Understanding the frequency, as the pitch of the instrument (in this case a tinkle bell instrument), the faster the wave repeats itself (given the frequency), the higher the pitch of the note.
-It is important to mention that any physical object or phenomenon that emits noise periodically (hundreds or thousands of times per second) must have a perceptible tone.
-However, not all sound is periodic.
-For example, since white noise represents static sound, it is just a uniform distribution of audible frequencies, or Gaussian distribution.
-Therefore, because white noise is non-periodic, it does not have a noticeable pitch.
+The visual representation of the objects were represented in three ways: the original light curves, the sonification waveforms and the spectrograms.
+All plots are aligned along the time axis for direct comparison.
+For each blazar, the waveforms represent the audio signal as amplitude versus time, where amplitude corresponds to the perceived loudness of each note.
+Since the datasets were normalized prior to sonification, the maximum amplitude in all waveforms is 1.
+Importantly, the waveforms do not encode pitch; they only show how the loudness of the sound varies over time.
+The spectrograms were generated from the sonified audio rather than from the original light curve data.
+They display the distribution of frequencies in the sonification as a function of time.
+The vertical color bar indicates the intensity of the signal in each time–frequency bin.
+Warmer colors (e.g., red or yellow) correspond to stronger signal components, while cooler colors (e.g., blue or black) indicate weaker ones.
+The presence of multiple bright frequencies within a single time bin arises both from the harmonic overtones characteristic of the tinkle bell instrument and from temporal overlap of successive notes.
+Together, the light curves, waveforms, and spectrograms provide complementary perspectives on the same data: the original flux variations, the time-domain structure of the sonification, and the distribution of its pitch components across time.
 4 Discussion
 Using a bell as a sonification instrument in this part of the production enhances clarity and auditory distinction.
-Its resonant tone makes it easier to perceive important events in the data without blending with other sounds.
+Its resonant tone, rich in overtones, has been shown to facilitate the auditory salience of events by making them stand out from the background (Kramer et al., 1999; Grond & Berger, 2011).
+Previous studies in auditory display and sonification highlight that instruments with clear attack and sustained resonance improve the perceptual separation of events, reducing masking effects and allowing listeners to identify patterns more easily (Hermann, 2008).
+Therefore, the tinkle bell timbre was chosen to ensure that important variations in the data could be perceived without blending with other sounds.
 Additionally, its sharp attack (well-defined onset) allows for precise marking of specific points in time, which is particularly useful for sonifying time series or discrete events.
-The bell also generates harmonics that enrich the tonal representation of data variations, helping to distinguish between different values or categories.
-For example, in the light curves presented here, higher flux values translate into a more intense or higher-amplitude sound that gradually fades over time, fully covering the gaps in the curve.
-While these changes are evident in waveforms, they may go unnoticed in light curves.
-However, the ear can detect abrupt shifts in intensity or volume.
-This is the advantage of sonification: what escapes the eye can be perceived by the ear.
+The bell timbre is also characterized by a rich harmonic spectrum: when a note is struck, besides the fundamental frequency, several overtones of different amplitudes are produced.
+These overtones increase the perceptual brightness and richness of the sound, which improves the listener’s ability to discriminate subtle variations between consecutive notes (Zwicker & Fastl, 1999).
+In practice, this means that small changes in the mapped data values are more easily perceived as distinct auditory events, as the harmonics reinforce pitch perception and provide additional auditory cues beyond the fundamental tone.
+For example, in the light curves presented here, the amplitude of each note is directly tied to the flux value of the corresponding data point: higher flux corresponds to a higher amplitude.
+The note length is constant for all points and is chosen to ensure that each note sufficiently covers the temporal spacing between consecutive measurements, effectively bridging gaps in the light curve.
+This approach allows abrupt changes in intensity or volume to be perceived by ear, even if they are not visually evident in the waveform or the original light curve.
+Consequently, sudden increases or decreases in flux are more perceptible auditorily, providing an additional modality for identifying significant variations in the data.
+This is the advantage of sonification: features that may be difficult to perceive visually, such as rapid short-term flares or subtle oscillations in the flux, can be detected by the ear.
+When sonified, these variations correspond to sudden changes in pitch or amplitude, allowing listeners to perceive patterns and events that might escape visual inspection of the light curves.
 Synchronising the three plots (light curve, waveform, and spectrogram) along the time axis provides a multidimensional representation of the data.
@@
 This makes it possible to identify transient events such as peaks, dips, eclipses, explosions, or periodic variations.
-The waveform, on the other hand, captures the full temporal evolution of the signal with high amplitude resolution.
-It acts as a detailed zoom into the temporal variations that might not be visible in the light curve.
-This allows for the analysis of precise dynamics, such as rapid pulses, subtle oscillations, or specific shapes.
-It also helps to distinguish noise from real signals and correlate light curve peaks with their exact amplitude and timing.
-The spectrogram decomposes the signal into its frequency components, showing how they evolve over time.
-It helps to identify events linked to specific frequency bands, such as peaks or oscillation modes and distinguishes noise, characterised by specific spectral patterns, from significant signals.
+The waveform captures the full temporal evolution of the signal, providing a zoom into temporal variations that might not be visible in the light curve.
+For example, the oscillations in Mrk 501 that appear as small irregularities in the light curve allow us to separate short-term variations from instrumental or observational noise.
+Similarly, the sonification and waveform visualization make it possible to correlate the peaks in the light curve with their amplitude and timing, such as the 224-day periodicity observed in Mrk 501 (Magallanes-Guijón, 2020), highlighting both the intensity and the temporal spacing.
+The spectrogram is computed from the sonified audio signal rather than from the original astronomical data.
+It decomposes this audio waveform into its frequency components, showing how the pitch content generated by the sonification evolves over time.
+For example, in the case of PKS J2134-0153 (Figure 7), short-lived X-ray and $\gamma$-ray flares that appear only as subtle bumps in the light curve emerge as transient bright bands in the spectrogram.
+Because the tinkle bell instrument produces a fundamental pitch together with inharmonic overtones, each mapped data event spreads across multiple frequency bands, which become visible in the spectrogram.
+This visualization therefore reveals the temporal distribution, harmonic structure, and overlap of sonified events, complementing both the light curve and the auditory perception.
 Combining these three plots provides a multiscale analysis:
   • The light curve offers a general view of the signal’s behaviour over time.
+For example, in OJ 287, the quasi-periodic variability is visible is visible in the light curve as a series of recurrent peaks, establishing the large-scale context of variability.
   • The waveform adds fine detail about amplitude variations.
+Rapid flares that appear only as small bumps in the light curve become clearly distinguishable in the waveform, allowing us to separate genuine short-term variability from instrumental noise.
   • The spectrogram reveals frequency properties, uncovering patterns not evident in the other two.
+In the case of Mrk 421, the use of the tinkle bell instrument generates harmonics that appear as multiple bright bands in the spectrogram.
+While this complicates the auditory perception, it enhances the visual identification of event timing and harmonic structure, making subtle features—such as overlapping flares—more apparent.
 This integrated approach enables the correlation of complex events.
@@
 This is relevant for identifying specific physical phenomena, such as periodic pulses, spectral bursts, or emission changes, often observed in the study of blazars or binary stars.
+The corresponding spectrogram of Mrk 501, for example, shows transient bright frequency optical band, indicating the presence of oscillations.
 Moreover, this combination is key to differentiate real signals from noise.
@@
 For transient events or explosions, the light curve can measure the event’s total duration and peak intensity, the waveform provides details about the pulse dynamics, and the spectrogram highlights the specific frequency bands involved.
-For example, in the observation of a blazar, the light curve tracks global brightness variations, the waveform reveals rapid underlying pulses invisible on a larger scale, and the spectrogram identifies harmonic components or specific spectral features.
+For instance, in case of OJ 049 a small fluctuation in the light curve initially appears as a possible flare.
+However, inspection of the waveform shows no coherent pulse dynamics, and the spectrogram lacks any associated frequency bands, confirming that the fluctuation is consistent with observational noise.
+The light curve reveals the overall flare duration and peak intensity, the waveform captures the sharp rise and decay dynamics, and the spectrogram highlights transient frequency components produced by the sonified signal.
+This integrated analysis of OJ 049 not only confirms which apparent flares are artifacts of noise but also reveals genuine variability signatures.
+In particular, during epochs of increased activity, the waveform displays brief amplitude surges that correspond to microflares lasting only a few hours, while the spectrogram shows simultaneous enhancements across multiple frequency bands.
+These correlated features indicate that the variability arises from real physical changes-most likely turbulent processes or shock propagation within the jet-rather than from instrumental or sampling effects.
+Thus, combining the three representations enables a clearer discrimination between stochastic noise and intrinsic variability in the blazar’s emission.
 The synchronised combination of light curve, waveform, and spectrogram provides a comprehensive and integrated view of the data: intensity, temporal dynamics, and frequency structure.
 This enhanced analysis, facilitates the identification of complex events and improves the interpretation of physical phenomena.
-As mentioned in the Introduction, there are already several works on data sonification in astronomy.
-These works provide important information from the perspective of sonifying the image, since the areas with the highest light intensity can be distinguished with the ear.
-In addition, they have a high pedagogical value, since they allow the public to have a didactic and playful approach to astronomy.
-However, these works do not show on the time line how are the variations, trends, cyclicity, periodicity, seasonality of the phenomena or of the radiative processes.
-In this sense, the route that we have taken is that of sonification of the time series, or light curves, and the exercise that we have carried out has been on nine active galaxies that contain a supermassive black hole in their nucleus, an accretion disk that surrounds it, and a relativistic jet, so the radiation emission from these objects is multifrequency.
-Thus, sonification data for the Mrk 501, Mrk 1501, Mrk 421, BL Lacerta, AO 0235+164, 3C 66A (PKS 0219+428), OJ 049 (PKS 0829+046), OJ 287, and PKS J2134-0153 blazars were presented with their associated light curves, waveforms and spectrograms.
+In the case of OJ 049, the joint examination of the three representations reveals that apparent fluctuations in the optical light curve without corresponding features in the waveform or spectrogram are consistent with observational noise.
+Conversely, when a true flare occurs, the waveform shows coherent amplitude rises and decays, and the spectrogram displays transient frequency bands linked to the event.
+Similarly, in Mrk 421, correlated peaks across both optical and X-ray spectrograms highlight multiwavelength variability episodes, confirming that the combined visual–auditory approach enhances the detection and understanding of real astrophysical processes.
+As mentioned in the Introduction, there are already several works on data sonification in astronomy, which explore different mapping strategies and applications.
+For instance, Dubus & Bresin (2013) provided a systematic review of sonification mapping strategies for physical quantities, establishing a foundation for perceptual and analytical approaches.
+More recent studies, such as Guiotto Nai Fovino et al.
+(2024), evaluated the effectiveness of sonification for time-series data exploration, highlighting its potential to enhance data comprehension beyond visual analysis.
+Likewise, Tucker Brown et al.
+(2022b) tested the efficacy of sonification for signal detection in light curves through the ASTRONIFY framework, demonstrating measurable improvements in identifying transient phenomena.
+From a broader perspective, Mis examined how sound experts perceive astronomy sonification projects, emphasizing both their scientific and communicative dimensions.
+Similarly, Casado & García (2024) discussed multimodal analysis approaches, integrating auditory and visual representations to strengthen the interpretative process of astrophysical data.
+These works not only confirm the analytical value of sonification but also underscore its pedagogical potential (Zanella et al., 2022), as it enables the public to experience complex astronomical data in an engaging manner.
+While previous works on sonification have focused primarily on mapping static data or images, they generally do not display how variations, trends, cyclicity, periodicity, or seasonality evolve over time.
+In contrast, our approach focuses on the sonification of temporal data – specifically, light curves – to capture the dynamic behavior of astrophysical phenomena.
+We applied this method to nine active galaxies hosting supermassive black holes with surrounding accretion disks and relativistic jets, whose radiation exhibits multifrequency variability.
+The sources analyzed include Mrk 501, Mrk 1501, Mrk 421, BL Lacertae, AO 0235+164, 3C 66A (PKS 0219+428), OJ 049 (PKS 0829+046), OJ 287, and PKS J2134-0153.
+For each case, we present the sonified data alongside their corresponding light curves, waveforms, and spectrograms, allowing for a multiscale and multimodal interpretation of their emission behavior.
 5 Conclusions
@@
 The integration of these methods not only enables a richer interpretation of the data but also provides perspectives for scientific communication, particularly for visually impaired communities.
-Although this approach does not aim to definitively demonstrate periodicities, the auditory representation offers an innovative way to identify potential patterns, correlate events across different time scales, and explore the complex dynamics of these astrophysical objects.
-This work lays the groundwork for future research to delve deeper into the potential of sonification to highlight regularities and transient phenomena in astronomical data.
-This multimodal approach contributes to a more inclusive and diverse understanding of blazar behaviour, paving the way to expand scientific analysis through methods that integrate both visual and auditory perception.
-The radiation processes in the studied galaxies are multifrequency, so the data we have processed for sonification correspond to radio, optical, X-ray, and gamma-ray bands.
-Thus, the auditor can listen to the recorded signals that are emitted due to the radiative processes that distinguish the components of the active nucleus of the galaxy.
-Data were obtained from seven observatories: UMRAS, AAVSO, Astrogeo-VLBI, Swift, Fermi, and ZTF.
-The graphs presented for each of the blazars are specially useful for detecting changes over time, as well as getting a closer idea of sonification in visual terms.
-In these one can visualise what has been sounded, the changes in intensity (volume) and frequency (tone).
-It is important to say that other fields in astronomy in which the sonification of data is present are in matters of education and public participation.
-There is a record of how sonifying images is impacting the dissemination of astronomy, and how citizen participation is increasing (Zanella et al., 2022).
-The sonification of data is a way with which substantial information can be extracted, such as the search for periodicities, power increases, regularities, absences of information, etc.
-In addition to visualising the data, we have sonified it with the purpose of disseminating our advances to an even broader spectrum of society, at the same time that we investigate auditory events through the sense of hearing that help to distinguish periodicities.
-Thus, the central interest of this work has been to disseminate the analysis of astrophysical data from a technique that is attractive to science in general, which is listening to the phenomena.
-In this sense, this is a work to bring the general public closer to astrophysics, from the sonification of the light curves of the active nuclei of galaxies, but also from the concern to socially include in science sectors of the population are blind or vision impaired (BVI) (Pérez-Montero, 2019).
+In addition, these sonifications serve a dual purpose for the visually impaired community.
+First, they provide a means for BVI scientists to perceive patterns, periodicities, and transient events directly through auditory cues, supporting independent data exploration.
+Second, they can be used as an educational and communicative tool, enabling the broader BVI audience to access and understand complex astrophysical phenomena.
+For example, short-lived oscillations that appear as subtle fluctuations in a light curve can become more perceptually salient in the sonification, allowing listeners to detect features that are otherwise difficult to visualise.
+Although this study does not aim to definitively demonstrate periodicities, the auditory representation provides an approach to detect potential patterns, correlate events across different time scales, and explore the complex dynamics of these astrophysical objects.
+By demonstrating how sonification can complement traditional visualizations, this work establishes a methodological foundation for future research, including systematic searches for periodicities, transient events, and multiwavelength correlations in astronomical data.
+This multimodal approach enables a more inclusive and diverse understanding of blazar behaviour by allowing data to be explored both visually and auditorily.
+For example, visually impaired (BVI) scientists can access the temporal dynamics, amplitude variations, and frequency patterns in light curves through sonification, while sighted researchers can combine waveforms and spectrograms with sound to identify subtle features that might be missed visually.
+This integration enhances accessibility and provides complementary perspectives on the same data, improving the detection of flares, oscillations, and other transient phenomena.
+The radiation processes in the studied blazars are multi-frequency, including synchrotron emission, inverse Compton scattering, and thermal radiation from the accretion disk.
+These processes produce variability in different energy bands (radio, optical, X-ray, and gamma-ray), which we sonified.
+As a result, the listener can perceive changes associated with specific radiative mechanisms: for example, rapid optical flares generated by synchrotron emission correspond to short, high-pitched notes, while slower variations in gamma-ray flux appear as longer, lower-pitched tones.
+In this way, the sonification provides an auditory representation of the physical processes driving the emission, linking observable features to their underlying astrophysical origins.
+The graphs presented for each blazar are particularly useful for understanding the sonification in visual terms.
+By pairing the visualizations with the corresponding sounds, one can see how changes in flux or luminosity are reflected in the amplitude (volume) and frequency (pitch) of the sonified signal.
+This visual-auditory correspondence aids in interpreting complex temporal patterns, verifying features detected by ear, and providing a more comprehensive understanding of the data.
+The sonification of data allows the extraction of substantial information, such as searching for periodicities, detecting increases in power, identifying regularities, or noting absences of data.
+Beyond visualizing the data, we sonified it to extend the accessibility of our results to a broader audience.
+In particular, this includes both the general public, who can gain an engaging and intuitive understanding of astrophysical phenomena through sound, and blind or visually impaired (BVI) (Pérez-Montero, 2019).
+individuals, including both laypeople and professional astronomers.
+By translating complex temporal patterns into auditory cues, the sonifications provide a pathway for BVI audiences to perceive and interpret features that might otherwise rely solely on visual inspection.
 Acknowledgements
@@
 The data underlying this article are available in the article and in its online supplementary material, together with the complementary webpage: https://www.guijongustavo.org/datasonification.
+Appendix A Blazar Sample and Specifications
+For each dataset, the time resolution and integration per data point are determined by the corresponding instrument or database.
+In the optical band, ZTF data consist of single exposures of 30 s (Masci et al., 2019), while AAVSO records typically correspond to integrations of 60–300 s depending on the observer (Kafka & Beck, 2021).
+In the radio band, UMRAO data points were obtained from integrations of approximately 30 minutes per observation (Aller et al., 1985), and Astrogeo VLBI measurements correspond to individual interferometric sessions (Petrov et al., 2021).
+For the X-ray band, Swift-XRT provides observations with effective exposures ranging from a few hundred to a few thousand seconds (Burrows et al., 2005).
+In the case of the $\gamma$-ray band, Fermi-LAT light curves are generally binned into one-week intervals to ensure sufficient photon statistics (Ballet et al., 2023).
+These values represent the effective time resolution of the light curves used in this work.
+Each blazar selected had the following specifications:
+(1) Mrk 501, has a redshift $z$ = 0.034 that corresponds to a distance $\sim$ 140 Mpc and R.
+A.
+= 16h 53m 52.2s, Dec.
+= +39° 45’ 37”.
+It has been monitored in radio (Richards et al., 2011), optical (Smith et al., 2009), X-rays (Abdo et al., 2011), and $\gamma$-rays (Dorner et al., 2017).
+This object has several periodicities reported (Bhatta, 2019).
+Recently Magallanes-Guijón & Mendoza (2024) have reported a periodicity $\sim$ 224 days.
+The light curve in optical was built using the database from 1998 June 24 to 2021 September 12.
+The Swift database is from 2008 October 2 to 2021 October 8.
+And the $\gamma$-rays, covers a time interval from 2008 August 04 up to 2020 March 13.
+Table 2 shows specific information about the records and the light curve is presented in Figure 1.
+(2) Mrk 1501 is a FSRQ with redshift of $z$ = 0.089338 at a distance of $\sim$ 377 Mpc (Sargent & Searle, 1970).
+This blazar presents short-lived $\gamma$-ray emission flares Arsioli & Polenta (2018).
+It is in the Pisces’s constellation, with R.
+A.
+= 00h 10m, and Dec.
+= +10° 58’.
+(Markarian et al., 1989).
+For the optical, the dataset was taken from the ZTF database, from 2010 July 25 to 2021 November 25.
+In the case of X-rays the database comes from the Swift observatory from 2008 October 2 to 2021 March 22.
+For the period of 2008 August 4 to 2022 October 22, the dataset is from the Fermi database.
+Table 2 shows the summary of the dataset, and Figure 2 presents the light curves.
+(3) BL Lacerta was discovered in 1929 by Hoffmeister (1929) and Oke & Gunn (1974) measured its redshift $z$ = 0.07 which corresponds to a distance of $\sim$ 0.276 Gpc.
+It is in the constellation Lacertae, with R.
+A.
+= 22h 02m 43.3s, Dec.
+= +42° 16’ 40” (Miller et al., 1978).
+The dataset in radio for this object was obtained from UMRAO from 2009 September 1 to 2012 May 21.
+For optical, the dataset used was from AAVSO in an interim of 1969 December 14 to 2022 March 13.
+For X-rays, the time interval is 2008 October 2 to 2022 November 10, taken from the Swift Observatory.
+And for $\gamma$-rays, the time of observation was from 2008 August 4 to 2022 March 11.
+Table 2 presents the summaries of the dataset and the light curves are shown in Figure 3.
+[…]
+For AO 0235+164, the optical dataset was obtained from 1998 August 25 to 2020 September 13, and the database used was from AAVSO.
+For X-rays, the dataset used was from Swift from 2008 October 2 to 2017 December 7.
+The Fermi $\gamma$-rays database was taken from 2008 August 4 to 2020 April 5.
+Table 2 shows a summary of these details and Figure 4 presents the corresponding light curves.
+(5) 3C 66A has a redshift of $z$ = 0.34 corresponding to $\sim$ 1.4 Gpc (Torres-Zafra et al., 2018).
+This BL Lac object has been observed on optical wavelengths (e.
+g.
+Böttcher et al., 2005; Lainela et al., 1999).
+It has R.A.
+= 2h 23m 12s and Dec.= 43° 0.’7 reported by Errando et al.
+(2009) in $\gamma$-rays.
+The radio records from the Astro-geo VLBI were taken from 1997 January 11 to 2020 June 4.
+For optical, the dataset used was from AAVO from 1994 February 1 to 2022 February 10.
+The Swift X-rays database used is from 2008 October 2 to 2022 December 9.
+All these specifications are presented in Table 2 and the light curves are shown in Figure 5.
+(6) OJ 049 has a redshift $z$ = 0.18 $\sim$, corresponding to a distance of $\sim$ 760 Mpc (Falomo, 1991).
+It was detected as an X-rays source (Della Ceca et al., 1990) and as a GeV $\gamma$-rays source (Fichtel, 1994; von Montigny et al., 1995; Mattox, 1998; Mukherjee et al., 1997).
+Its R.
+A.
+= 08h 31m 48.8s and Dec.
+= $04^{\circ}$ 29’ 39” (Fiorucci, M.
+& Tosti, G., 1996).
+For this blazar the databases were AAVSO, Swift and Fermi.
+The corresponding time periods were from 2008 January 9 to 2020 June 23, 2002 October 9 to 2020 March 31, 2006 October 23 to 2021 December 31 and 2008 August 5 to 2020 April 6, respectively.
+Table 2 shows all these details and their corresponding light curves are presented in Figure 6.
+(7) PKS J2134-0153 has a redshift $z$ = 1.285, corresponding to a distance $\sim$ 5 Gpc (Truebenbach & Darling, 2017) with R.
+A.
+= 21h 34m 10.3095s and Dec.
+= $01^{\circ}$ 53’ 17.238” (Fomalont et al., 2000).
+For optical, the database is from ZTF with interim from 2009 October 4 to 2010 August 14.
+In X-rays, the dataset is from Swift from 2008 October 2 to 2014 October 8.
+For $\gamma$-rays, the database is Fermi, with records from 2008 August 4 to 2022 October 22.
+All these details are presented in Table 2.
+Figure 7 shows the corresponding light curves.
+(8) Mrk 421 has a R.
+A.
+= 11h 04m 27 31s and Dec.
+= $38^{\circ}$ 12’ 31.80”, located at a redshift $z$ = 0.031, which corresponds to a distance $\sim$ 130 Mpc) (Ulrich et al., 1975).
+This object shows rapid flux and polarisation variability (e.
+g., Fraija et al., 2017).
+For optical, the dataset is from AAVSO database, from 1984 June 22 to 2020 April 13.
+In the case of X-rays, the database used was Swift, from 2005 March 1 to 2020 April 1.
+For $\gamma$-rays, the database Fermi gave the dataset from 2008 June 5 to 2020 March 8.
+The summary of all this is presented in Table 2.
+The correspond light curves are shown in Figure 8.
+(9) OJ 287 is a BL Lacerta type quasar situated at a redshift of $z$ = 0.306 $\sim$ 4 Gly $\sim$ 1 Gpc with R.
+A.
+= 08h 54m 48.87s and Dec.
+= +20° 06’ 30.6”.
+This blazar is a candidate to host a binary supermassive black hole (Valtonen et al., 2006) with a periodicity of 12 years in its optical wavelength (Shi et al., 2007).
+For optical, the database was obtained from AAVSO, from 1987 September 26 to 2020 June 3.
+Swift provided the dataset of X-rays from 2008 November 12 to 2020 April 28.
+In the case of $\gamma$-rays the dataset is from Fermi, from 2008 April 4 to 2020 June 6.
+Table 2 show the records of this blazar.
+Figure 9 presents the corresponding light curves.
+Appendix B Tables
+In Tables 1 and 2, we use the following conventions: a Records refers to a single photometric or flux measurement in a given band, typically corresponding to one integration by the instrument or observer.
+The term Interim [d] indicates the total time span covered by the dataset, expressed as years, months, and days [y, m, d] between the first and last observation date, regardless of gaps.
+The column Total [d] refers to the number of distinct calendar days within that interval for which at least one measurement was obtained.
+Blazar R.A. Dec. Redshift Distance
+Mrk 501 16h53m52.2s +39°45’37” 0.034 140 Mpc
+Mrk 1501 00h10m +10°58’ 0.089338 377 Mpc
+BL Lacerta 22h02m43.3s +42°16’40” 0.07 0.276 Gpc
+AO 0235+164 02h35m52.6s +16°24’05” 0.940 3 Gpc
+3C 66A 02h23m12s +43°00’07” 0.34 1.4 Gpc
+OJ 049 08h31m48.8s +04°29’39” 0.18 760 Mpc
+PKS J2134-0153 21h34m10.3095s -01°53’17.238” 1.285 5 Gpc
+Mrk 421 11h04m27.31s +38°12’31.8” 0.031 130 Mpc
+OJ 287 08h54m48.87s +20°06’30.6” 0.306 1 Gpc
+Table 1: Summary of the nine blazars analyzed in this work, including equatorial coordinates (R.A., Dec.), redshift, and approximate distance.
+Blazar Band Records Interim Total Dates
+[y, m, d] [d] [d]
+Mrk 501 Optical 11,849 23, 2, 17 8,441 1998-06-24 – 2021-09-12
+X-rays 28,000 12, 7, 9 4,607 2008-10-02 – 2021-10-08
+$\gamma$-rays 4,199 11, 7, 9 4,239 2008-08-04 – 2020-03-13
+Mrk 1501 Optical 1,255 11, 4, 0 4,141 2010-07-25 – 2021-11-25
+X-rays 11,150 12, 5, 20 4,554 2008-10-02 – 2021-03-22
+$\gamma$-rays 5,146 14, 2, 18 5,192 2008-08-04 – 2022-10-22
+BL Lacertae Radio 413 12, 5, 25 4,560 2009-09-01 – 2012-05-21
+Optical 10,117 52, 2, 27 19,082 1969-12-14 – 2022-03-13
+X-rays 100,418 14, 1, 8 5,152 2008-10-02 – 2022-11-10
+$\gamma$-rays 4,892 13, 7, 7 4,967 2008-08-04 – 2022-03-11
+AO 0235+164 Optical 196 22, 9, 17 8,055 1998-08-25 – 2020-09-13
+X-rays 61,146 9, 2, 5 3,353 2008-10-02 – 2017-12-07
+$\gamma$-rays 4,223 11, 8, 1 4,262 2008-08-04 – 2020-04-05
+3C 66A Radio 101 23, 4, 22 8,545 1997-01-11 – 2020-06-04
+Optical 8,470 28, 0, 9 10,236 1994-02-01 – 2022-02-10
+X-rays 3,163 14, 2, 7 5,181 2008-10-02 – 2022-12-09
+$\gamma$-rays 4,223 11, 8, 1 4,262
+OJ 049 Optical 76 17, 5, 22 6,383 2008-01-09 – 2020-06-23
+X-rays 16,111 15, 2, 8 5,548 2002-10-09 – 2020-03-31
+$\gamma$-rays 4,196 11, 8, 1 4,262 2008-08-05 – 2020-04-06
+PKS J2134-0153 Optical 73 0, 10, 10 314 2009-10-04 – 2010-08-14
+X-rays 4,123 6, 0, 6 2,197 2008-10-02 – 2014-10-08
+$\gamma$-rays 4,223 11, 8, 1 4,262 2008-08-04 – 2022-10-22
+Mrk 421 Optical 18,167 35, 9, 20 13,079 1984-06-22 – 2020-04-13
+X-rays 160,346 15, 1, 0 5,510 2005-03-01 – 2020-04-01
+$\gamma$-rays 4,129 11, 9, 3 4,294 2008-06-05 – 2020-03-08
+OJ 287 Optical 13,754 32, 8, 6 11,939 1987-09-26 – 2020-06-03
+X-rays 9,464 11, 5, 16 4,185 2008-11-12 – 2020-04-28
+$\gamma$-rays 4,410 12, 2, 2 4,446 2008-04-04 – 2020-06-06
+Table 2: Summary of all electromagnetic bands studied for the nine blazars.
+Records corresponds to a single flux measurement.
+Interim [d] is the total time span covered by the observations, expressed in years, months, and days [y, m, d].
+Total [d] indicates the number of distinct days within the interim for which at least one measurement is available.